ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 20C15, 20C33 Darafsheh, M. R., University of Tehran, Iran The character table of the group GL_{2}(q) when extended by a certain group of order two Let G denote either of the groups GL_{2}(q) or SL_{2}(1q). Then θ : G → G given by θ(A) = (A^{−1})^{t}, where A^{t} denotes the transpose of the matrix A, is an automorphism of G. Therefore we may form the group G.⟨θ⟩ which is the split extension of the group G by the cyclic group ⟨θ⟩ of order 2. Our aim in this paper is to find the complex irreducible character table of G.⟨θ⟩.